OpenStudy (faiqraees):

A crystal lattice has 27 atoms. Each atom has a mass of 3.5x10^-25 kg. The density of the crystal is 9.2x10^3kgm^-3. What is the shortest distance between the centres of two adjacent atoms? @phi

OpenStudy (faiqraees):

@raffle_snaffle

OpenStudy (faiqraees):

V= 8x³ D = M/V V= M/D V = 27M/D 8x³= 27x3.5x10^-25 /9.2x10^3 What's wrong here?

OpenStudy (raffle_snaffle):

what equation are you using?

OpenStudy (faiqraees):

I mentioned the equations

OpenStudy (faiqraees):

Should I explain how I got the equations?

OpenStudy (raffle_snaffle):

what is Volume of a sphere? Equation?

OpenStudy (faiqraees):

Volume of a lattice its a cube

OpenStudy (faiqraees):

Volume = length³

OpenStudy (raffle_snaffle):

but it wants to know the distance between the atoms.

OpenStudy (faiqraees):

yeah so we will then divide the length by 2

OpenStudy (raffle_snaffle):

we don't know the dimension of the cube or where the atoms are inside the cube.

OpenStudy (raffle_snaffle):

we are relying on the atoms to tell us where they orbit inside the cube.

OpenStudy (faiqraees):